# Multi-Physics-Enhanced Bayesian Inverse Analysis: Information Gain from Additional Fields

### v1

@@ 4.2.2 Observations
We generated the synthetic observations based on an evaluation of the three-field porous medium model \$M(Eˆh, Eˆd)\$ with the ground-truth parameters \$Eˆh_true= $6\text{\,}\mathrm{k}\mathrm{P}\mathrm{a}$\$ and \$Eˆd_true= $12\text{\,}\mathrm{k}\mathrm{P}\mathrm{a}$\$. Starting the simulation from end-expiration, all observations were obtained after $2.5\text{\,}\mathrm{s}$ at end-inspiration. We extracted the displacements and blood volume fractions at the observation locations shown in Figure 12. For the displacements, we used \$N_obs, 1 = 7\$ observations and for the blood volume fractions, we used \$N_obs, 2 = 2\$ observations. Note that each displacement observation is a vector with a component in \$c_1\$-direction and a component in \$c_2\$-direction. To obtain the final observations, we added zero-mean Gaussian noise sampled from a Sobol’ sequence [40] according to Equation 3. The variance of the noise \$σ_jˆ2\$ was derived from the signal-to-noise ratio \$SNR_j\$ of each field \$j\$ (see Equation 12).
@@
4.2.3 Probabilistic model & inference method
For the prior distribution of the parameters of interest \$Eˆh\$ and \$Eˆd\$, we used a truncated normal distribution according to Table 3. The observations were incorporated with the Gaussian multi-physics log-likelihood function presented in Equation 4. The likelihood variance \$σ_jˆ2\$ of each field \$j\$ was always identical to the variance used to sample the observational noise as described in Section 4.2.2. With this prior and likelihood, we evaluated the posterior on a \$50×50\$ grid of \$Eˆh\$ and \$Eˆd\$. The grid points in each dimension were evenly spaced with respect to the cumulative distribution function of the prior. The model evaluations for the grid-based likelihood evaluations were scheduled in parallel using our open-source solver-independent multi-query framework QUEENS [43].
@@ 4.2.5 Relative increase in information gain for varying coupling strength and signal-to-noise ratios
{tikzpicture}
@@
[…]
Figure 14: The relative increase in information gain (RIIG) for \$k_J=-1.5\$ (top left), \$SNR_1=10\$ (top center), and \$SNR_2=10ˆ6\$ (top right) while varying the two remaining parameters. We also combine these three plots in a single, three-dimensional plot (bottom). The results demonstrate that for the example at hand, the RIIG from the single-physics to the multi-physics posterior is particularly high when the first field has a low SNR, the second field has a high SNR, and the coupling between the fields is strong 6. Even for weak coupling (\$k_J ∈[-0.5, -0.1]\$), the RIIG is still notable if the \$SNR_1\$ is sufficiently low and the \$SNR_2\$ sufficiently high.
@@
% Method 2nd figure In addition to the single multi-physics posterior and its information gain in Figure 13, we also analyze the RIIG for a range of first-field and second-field observations as well as for different coupling parameters \$k_J\$. While we varied the number and SNR of the second-field observations in section 4.1, we now vary the SNRs of both the first-field and second-field observations and the coupling parameter \$k_J\$. We evaluated the resulting RIIG on a \$50×50×10\$ grid of \$SNR_1\$, \$SNR_2\$, and \$k_J\$. The grid points of the SNRs were log-spaced, and the grid points of \$k_J\$ were evenly spaced.
@@
% Results & Discussion 2nd figure Figure 14 shows the resulting RIIG when varying these three parameters. The results demonstrate that the RIIG is higher for a lower SNR in the first field and a higher SNR in the second field. This is to be expected since a lower SNR in the first field means that observations of this field are less informative, resulting in an increasingly uninformative single-physics posterior. As a result, the single-physics information gain decreases, which increases the RIIG. Conversely, a higher SNR in the second field means that observations of this field are more informative. This increases the multi-physics information gain, which in turn increases the RIIG. Figure 14 further illustrates that the RIIG is higher the stronger the coupling between the two fields. This observation also aligns with expectations since a stronger coupling implies that the second field is influenced more strongly by the first field, which depends on the parameters of interest. We also observe that the RIIG is zero when the coupling parameter \$k_J\$ is zero. A coupling parameter of zero means that the second field does not depend on the first field. Hence, the second field contains no information about the parameters of interest, and the RIIG is zero. Furthermore, the RIIG is still notable even for weak coupling between the fields (\$k_J ∈[-0.5, -0.1]\$) if the \$SNR_1\$ is sufficiently low and the \$SNR_2\$ sufficiently high. In cases where the \$SNR_2\$ is not high enough, we assume that a sufficiently large number of second-field observations can compensate for this lack of information and increase the RIIG as shown in section 4.1. Our findings indicate that although strong coupling yields a higher RIIG, observations of a weakly coupled second physical field can still enhance a Bayesian inverse analysis.
@@ 5 Conclusion
This work proposes the approach of multi-physics-enhanced Bayesian inverse analysis. We demonstrate that this approach can substantially reduce the uncertainty in a previously uninformative posterior distribution by adding observations of additional physical fields. This reduction in uncertainty is quantified by evaluating the information gain from the prior to the posterior distribution. To assess the effect of the multi-physics observations, we compared the information gain with and without observations from a second physical field. Our results show that even a few or noisy additional observations of a second physical field can reduce uncertainty considerably. Even additional observations from just a one-way coupled or weakly coupled physical field resulted in a notable uncertainty reduction. Overall, our findings highlight the largely neglected potential of multi-physics observations to enhance Bayesian inverse analysis. This considerable potential can be leveraged with minimal effort in the frequent case where multi-physics observations are easily obtainable and a corresponding multi-physics model is available or easy to implement. This implementation effort is particularly low for one-way coupled or uncoupled multi-physics models. Building on our findings, we encourage experimenters, modelers in academia and industry, and developers of inverse methods to leverage the potential of additional physical fields and address previously intractable inverse problems.

### v2

@@ 4.2.2 Observations
We generated the synthetic observations based on an evaluation of the three-field porous medium model \$M(Eˆh, Eˆd)\$ with the ground-truth parameters \$Eˆh_true= $6\text{\,}\mathrm{k}\mathrm{P}\mathrm{a}$\$ and \$Eˆd_true= $12\text{\,}\mathrm{k}\mathrm{P}\mathrm{a}$\$. Starting the simulation from end-expiration, all observations were obtained after $2.5\text{\,}\mathrm{s}$ at end-inspiration. We extracted the displacements and blood volume fractions at the observation locations shown in Figure 12. For the displacements, we used \$N_obs, 1 = 7\$ observations and for the blood volume fractions, we used \$N_obs, 2 = 2\$ observations. Note that each displacement observation is a vector with a component in \$c_1\$-direction and a component in \$c_2\$-direction. To obtain the final observations, we added zero-mean Gaussian noise sampled from a Sobol’ sequence [undefam] according to Equation 3. The variance of the noise \$σ_jˆ2\$ was derived from the signal-to-noise ratio \$SNR_j\$ of each field \$j\$ (see Equation 12).
4.2.3 Probabilistic model & inference method
For the prior distribution of the parameters of interest \$Eˆh\$ and \$Eˆd\$, we used a truncated normal distribution according to Table 3. The observations were incorporated with the Gaussian multi-physics log-likelihood function presented in Equation 4, assuming i.i.d. Gaussian noise with known variances. The likelihood variance \$σ_jˆ2\$ of each field \$j\$ was always identical to the variance used to sample the observational noise as described in Section 4.2.2. With this prior and likelihood, we evaluated the posterior on a \$50×50\$ grid of \$Eˆh\$ and \$Eˆd\$. The grid points in each dimension were evenly spaced with respect to the cumulative distribution function of the prior. The model evaluations for the grid-based likelihood evaluations were scheduled in parallel using our open-source solver-independent multi-query framework QUEENS [undefap].
@@ 4.2.5 Relative increase in information gain for varying coupling strength and signal-to-noise ratios
[…]
[anchor=south west,inner sep=0] (img) at (0,0) Refer to caption;
@@
Figure 14: The relative increase in information gain (RIIG) for \$k_J=-1.5\$ (top left), \$SNR_1=10\$ (top center), and \$SNR_2=10ˆ6\$ (top right) while varying the two remaining parameters. Center: We also combine these three plots in a single, three-dimensional plot. For the example at hand, the RIIG from the single-physics to the multi-physics posterior is particularly high when the first field has a low SNR, the second field has a high SNR, and the coupling between the fields is strong 6. Even for weak coupling (\$k_J ∈[-0.5, -0.1]\$), the RIIG is still notable if the \$SNR_1\$ is sufficiently low and the \$SNR_2\$ sufficiently high. Bottom: The prior, likelihoods of each observed field \$y_1 = dˆt\$ and \$y_2 = εˆb\$, and posterior of the multi-physics-enhanced Bayesian inverse analysis (BIA) indicated in the center plot and already presented in Figure 13. The likelihood of the second field considerably constrains the posterior.
@@
% Method 2nd figure In addition to the single multi-physics posterior and its RIIG in Figure 13, we also analyze the RIIG for a range of first-field and second-field observations as well as for different coupling parameters \$k_J\$. While we varied the number and SNR of the second-field observations in section 4.1, we now vary the SNRs of both the first-field and second-field observations and the coupling parameter \$k_J\$. We evaluated the resulting RIIG on a \$50×50×10\$ grid of \$SNR_1\$, \$SNR_2\$, and \$k_J\$. The grid points of the SNRs were log-spaced, and the grid points of \$k_J\$ were evenly spaced.
@@
% Results & Discussion 2nd figure The top and center plots in Figure 14 show the resulting RIIG when varying these three parameters. Each point in these plots corresponds to a multi-physics-enhanced BIA. This is also illustrated by the bottom plots in Figure 14, which show the prior, the likelihoods of each observed field, and the multi-physics posterior at the indicated location in the center plot. This posterior was already presented in Figure 13. The likelihood of the additionally observed second field intersects with the likelihood of the initially observed first field only in a small region of the parameter space. This results in a concentrated posterior with corresponding RIIG of 0.68. The results presented in the top and center plots demonstrate that the RIIG is higher for a lower SNR in the first field and a higher SNR in the second field. This is to be expected since a lower SNR in the first field means that observations of this field are less informative, resulting in an increasingly uninformative single-physics posterior. As a result, the single-physics information gain decreases, which increases the RIIG. Conversely, a higher SNR in the second field means that observations of this field are more informative. This increases the multi-physics information gain, which in turn increases the RIIG. Figure 14 further illustrates that the RIIG is higher the stronger the coupling between the two fields. This observation also aligns with expectations since a stronger coupling implies that the second field is influenced more strongly by the first field, which depends on the parameters of interest. We also observe that the RIIG is zero when the coupling parameter \$k_J\$ is zero. A coupling parameter of zero means that the second field does not depend on the first field. Hence, the second field contains no information about the parameters of interest, and the RIIG is zero. Furthermore, the RIIG is still notable even for weak coupling between the fields (\$k_J ∈[-0.5, -0.1]\$) if the \$SNR_1\$ is sufficiently low and the \$SNR_2\$ sufficiently high. In cases where the \$SNR_2\$ is not high enough, we assume that a sufficiently large number of second-field observations can compensate for this lack of information and increase the RIIG as shown in section 4.1. Our findings indicate that although strong coupling yields a higher RIIG, observations of a weakly coupled second physical field can still enhance a Bayesian inverse analysis.
@@ 5 Conclusion
This work proposes a novel approach that we denote as multi-physics-enhanced Bayesian inverse analysis. This approach leverages observations of additional physical fields to identify parameters of single-physics computational models, avoiding the costly redesign of experimental setups to enhance single-physics observations. We demonstrate that the multi-physics-enhanced approach can substantially reduce the uncertainty in an uninformative posterior distribution. We quantify this reduction in uncertainty by evaluating the information gain from the prior to the posterior distribution. To assess the effect of the multi-physics observations, our work is the first to compare the information gain with and without observations from a second physical field. Our results show that even a few or noisy additional observations of a second physical field can reduce uncertainty considerably. Even additional observations from just a one-way coupled or weakly coupled physical field resulted in a notable uncertainty reduction. The uncertainty reduction through the second-field observations was generally higher the lower the signal-to-noise ratio (SNR) in the first field, the higher the SNR and the number of observations in the second field, and the stronger the coupling between the two fields. Overall, our findings highlight the largely neglected potential of multi-physics observations to enhance (Bayesian) inverse analysis. This considerable potential can be exploited with minimal effort in the frequent case where multi-physics observations are easily obtainable and a corresponding multi-physics model is available or easy to implement. This implementation effort is particularly low when the multi-physics model used for the inverse analysis has either one-way coupled or uncoupled physical fields. Building on our findings, we encourage experimenters, modelers in academia and industry, and developers of inverse methods to leverage the potential of additional physical fields and address previously intractable inverse problems.